# From Prompt Engineering to Prompt Science With Human in the Loop

### 01 Jan 2024 v1

@@ Abstract.
As LLMs make their way into many aspects of our lives, one place that warrants increased scrutiny with LLM usage is scientific research. Using LLMs for generating or analyzing data for research purposes is gaining popularity. But when such application is marred with ad-hoc decisions and engineering solutions, we need to be concerned about how it may affect that research, its findings, or any future works based on that research. We need a more scientific approach to using LLMs in our research. While there are several active efforts to support more systematic construction of prompts, they are often focused more on achieving desirable outcomes rather than producing replicable and generalizable knowledge with sufficient transparency, objectivity, or rigor. This article presents a new methodology inspired by codebook construction through qualitative methods to address that. Using humans in the loop and a multi-phase verification processes, this methodology lays a foundation for more systematic, objective, and trustworthy way of applying LLMs for analyzing data. Specifically, we show how a set of researchers can work through a rigorous process of labeling, deliberating, and documenting to remove subjectivity and bring transparency and replicability to prompt generation process. A set of experiments are presented to show how this methodology can be put in practice.
@@ 3.2. Codebook building as a way to construct prompts
4. Experiments
As a demonstration of the methodology explained in the previous section, we will now walk through its execution for a specific research application. The problem to address in this research is that of auditing an LLM. While there have been many recent efforts in this area (Mökander et al., 2023; Rastogi et al., 2023), we still lack systematic and scientifically rigorous approaches. We attempted to address this using a simple, but effective method – prompting the same question differently to the given LLM to see how differently it responds. A detailed discussion of the auditing problem for an LLM is beyond the scope for the present article. Instead, our focus here is on constructing the set of probes that could be used for that auditing task.
Let us start by identifying different components for the pipeline in Figure 4.
  • Data: Original set of questions that we could use to probe an LLM.
  • Responses: Different prompts generated using the original question. These prompts can be then given to the LLM being investigated.
  • Criteria: We need to make sure that the responses generated are still representing the original question, and yet different enough to be meaningful for probing an LLM. In other words, our main criteria here are relevance (to the original question) and diversity (among the generated responses/prompts).
  • Application: We will use the responses that meet our criteria of relevance and diversity as prompts to the LLM being investigated. The idea here is if the original question calls for factual responses and if the LLM being audited gives meaningfully different answers with that question asked differently or by different people/persona, then the said LLM need further probing as there may be issues of bias or hallucination.
To operationalize the method proposed in Section 3.2, we started with setting up the initial version of the pipeline. The initial prompt was set as “Generate five different questions for the following question that represent the same meaning, but are different.” for the LLM, we used Mistral 7-B and the data for original questions came from TruthfulQA dataset (Lin et al., 2021). Given the application – generating different versions of the same question for auditing an LLM – we set the criteria for responses generated by Mistral as relevance and diversity. With this, phase 1 was over and we moved to phase 2.
For phase 2, we used two researchers with a good understanding of the application and qualitative assessment. They were tasked with iteratively developing and finalizing the criteria for assessing the generated responses (different versions of the original question). The assessors started with a 7-point Likert scale for assessing relevance and a binary scale for assessing diversity. 10 original questions were sent through the LLM with the initial prompt to generate 10 sets of responses. We decided to limit each set to contain five variations. The two assessors labeled these 10 sets separately. They marked for ‘relevance’ for 50 questions and marked for ‘diversity’ for 10 sets.
Once done, we measured their agreement, which turned out to be very low. The problem seemed to stem from 7-levels of relevance scale and the two researchers’ understanding of what each level meant. This resolution for the scale did not seem necessary. We, therefore, decided to change the scale for relevance to three levels: low-medium-high. Another round of coding on 10 different input data was done. Once again, we had relevance assessment for 50 data points and diversity assessment for 10 data points.
We used Cohen’s kappa for measuring ICR, and found that the values for this measure for relevance was 0.4634 and for diversity was 0.6154. These are considered to be moderate levels of agreements (Ritchie et al., 2003). After a discussion between the two researchers, the definitions of relevance and diversity were slightly revised and the coding process was repeated with a new set of data. After one more iteration, we achieved 0.6418 and 0.8276 for Cohen’s kappa values for relevance and diversity, respectively. These are considered to be substantial levels of agreement (Ritchie et al., 2003). At this point, we decided to finish phase 2. The primary outcome was the codebook that contained detailed instructions about how to evaluate relevance and diversity in this particular context. An important byproduct was the enhanced understanding that the two researchers now possessed for assessing the outputs of the LLM in a reasonably objective and consistent manner.
Figure 5. Final prompt template.
We now proceeded with phase 3. The criteria for evaluating the outputs of the LLM was set. We needed to now focus on revising the prompt. Specifically, we wanted to have a prompt that consistently generates outputs (a set of questions) that could achieve high enough relevance and diversity for their use in a downstream application. We set the threshold for this to be 75%. Similar to the previous phase, the same two researchers marked the generated responses for relevance and diversity using a new set of input data. The initial round of assessment showed that only around 50% of the outcomes met the goal of relevance (medium or high) and diversity. To change this, the original prompt was modified to include more details about the application and the criteria. Then this process was repeated. After two more iterations, the LLM was generating outcomes that were relevant and diverse for at least 80% of the cases. At this point, the process for phase 3 was complete.
For phase 4, a senior researcher tested the whole pipeline with the final versions of the prompt and the criteria with a few rounds of random samplings. They made a few minor adjustments to the prompt for improved readability and generalizability. The final prompt is shown in Figure 5. At this point, we can be assured that this prompt will lead to reasonably consistent, reliable, and high quality outputs. The prompt can also be easily modified to suit small variations in the downstream applications without having to repeat the whole method. The documentations from the discussions and design choices during phases 2 and 3 could be shared, much like source code for an application, to aid in explainability and extensibility of this pipeline.
@@ 5. Conclusion
Creatively or iteratively designing prompts to produce desirabile outcomes, what is often referred to as prompt engineering, does not address these concerns. In fact, prompt engineering may even worsen the problems by focusing too much on bending the process to generating the desired outcomes than to developing a verifiable, validated, and replicable process. In this article, we showed how prompt engineering can be turned into prompt science by introducing a multi-phase process with human-in-the-loop. Specifically, we divided the prompt to application pipeline into two major phases that provide two levels of verification: one for the criteria for evaluating the responses generated by the LLM, and the other for the prompt for generating the responses. Through a method inspired by qualitative coding for codebook development, we showed how to have scientific rigor in developing a reliable prompt and getting trustworthy response for a downstream application. This method was demonstrated using a case study that used an LLM for generating questions for auditing another LLM.
@@
Acknowledgements.
The experiments described in Section 4 are results of hard work by Maryam Amirizaniani, Adrian Lavergne, Elizabeth Okada, and Jihan Yao. We are also thankful to Aman Chadha and Tanya Roosta for their valuable comments and guidance during this process.

### 06 Feb 2024 v2

@@ Abstract.
As LLMs make their way into many aspects of our lives, one place that warrants increased scrutiny with LLM usage is scientific research. Using LLMs for generating or analyzing data for research purposes is gaining popularity. But when such application is marred with ad-hoc decisions and engineering solutions, we need to be concerned about how it may affect that research, its findings, or any future works based on that research. We need a more scientific approach to using LLMs in our research. While there are several active efforts to support more systematic construction of prompts, they are often focused more on achieving desirable outcomes rather than producing replicable and generalizable knowledge with sufficient transparency, objectivity, or rigor. This article presents a new methodology inspired by codebook construction through qualitative methods to address that. Using humans in the loop and a multi-phase verification processes, this methodology lays a foundation for more systematic, objective, and trustworthy way of applying LLMs for analyzing data. Specifically, we show how a set of researchers can work through a rigorous process of labeling, deliberating, and documenting to remove subjectivity and bring transparency and replicability to prompt generation process.
@@ 4. Conclusion
Creatively or iteratively designing prompts to produce desirabile outcomes, what is often referred to as prompt engineering, does not address these concerns. In fact, prompt engineering may even worsen the problems by focusing too much on bending the process to generating the desired outcomes than to developing a verifiable, validated, and replicable process. In this article, we showed how prompt engineering can be turned into prompt science by introducing a multi-phase process with human-in-the-loop. Specifically, we divided the prompt to application pipeline into two major phases that provide two levels of verification: one for the criteria for evaluating the responses generated by the LLM, and the other for the prompt for generating the responses. Through a method inspired by qualitative coding for codebook development, we showed how to have scientific rigor in developing a reliable prompt and getting trustworthy response for a downstream application.